OpenStudy (scott208):

There are three different positive integers whose reciprocals add up to 1. What is the product of those three integers?

ganeshie8 (ganeshie8):

Tried anything?

OpenStudy (aaronandyson):

Reciprocals means this: 5/4 reciprocal is 4/5

OpenStudy (scott208):

well i mean at first i thought the three numbers were 3. But then i saw that they were all supposed to be different

OpenStudy (scott208):

and @aaron i know what a reciprocal is

ganeshie8 (ganeshie8):

As a start, setup an equation

OpenStudy (scott208):

1/z+1/x+1/x=1 x*x*x=?

OpenStudy (aaronandyson):

Let the 3 number be x,y,z respectively. So, \[\frac{ 1 }{ x } + \frac{ 1 }{ y } + \frac{ 1 }{ z }= 1\]

OpenStudy (aaronandyson):

Take the LCM \[\frac{ yz }{ xyz } + \frac{ xz }{ xyz } + \frac{ xy }{ xyz } =1\]

OpenStudy (scott208):

yeah that's where i'm stuck

OpenStudy (alekos):

product xyz = yz+xz+xy

OpenStudy (alekos):

the only saving grace is that they're integers

OpenStudy (aaronandyson):

xyz = yz+xz+xy Now we have to use the try and error method we have substitute x , y and z with number such that the sum of their reciprocal is 1.

OpenStudy (aaronandyson):

I think I found the 3 numbers there 2,3,and 6 1/2 + 1/3 + 1/6 LCM = 6 3/6 + 2/6 + 1/6 3+2+1 = 6 and 6/6 = 1

OpenStudy (alekos):

looks good, but are there any others?

OpenStudy (scott208):

is it 2,3, and 6? because 1/2=3/6 and 1/3=2/6 and 2/6+3/6+1/6=6/6=1?

OpenStudy (aaronandyson):

Yep! :)

ganeshie8 (ganeshie8):

Nice

OpenStudy (scott208):

thank you so much!